Circular No. 7638

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001ci IN NGC 3079 A. V. Filippenko and R. Chornock, University of California at Berkeley,

report that a CCD spectrum (range 390–1000 nm), obtained on May 30 UT with the Keck-II 10-m reflector, reveals that the ‘possible supernova’ in NGC 3079 (cf. IAUC 7618) is indeed a supernova, of type Ic, ∼ 2–3 weeks past maximum brightness. The spectral shape is similar to that of SN 1990U at 12 days past maximum (Matheson et al. 2001, A.J. 121, 1648), but extinguished by 5–6 mag at visual wavelengths.

NO SUPERNOVA 2001bh IN UGC 11635 Filippenko and Chornock add that a CCD spectrum (range 390–1000

nm), obtained on May 30 UT with the Keck-II 10-m reflector, reveals that SN 2001bh (cf. IAUC 7622) is a Galactic star at essentially zero redshift. The specific absorption lines suggest a mid-K to late-K spectral type, although the overall continuum shape is significantly bluer.

COMET C/2001 A2 (LINEAR) Ephemeris extension to IAUC 7600 (elements on MPEC 2001-K14):

1999TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
May 31 5^{h}00.^{m}71 −26^{o}58.^{′}8 0.485 0.789 49.^{o}4 102.^{o}8 4.4
June 2 4 52.14 −27 17.7 0.462 0.796 49.5 104.3 4.3
4 4 42.60 −27 29.1 0.439 0.805 49.9 105.4 4.3
6 4 32.00 −27 31.5 0.416 0.816 50.5 106.3 4.2
8 4 20.29 −27 23.5 0.395 0.828 51.4 106.8 4.2
10 4 07.41 −27 03.6 0.373 0.842 52.5 106.9 4.1
12 3 53.28 −26 29.8 0.353 0.857 54.0 106.5 4.1
14 3 37.87 −25 40.3 0.334 0.873 55.9 105.7 4.0
16 3 21.13 −24 32.8 0.315 0.891 58.2 104.2 4.0
18 3 03.04 −23 05.2 0.299 0.910 61.1 102.2 4.0
20 2 43.65 −21 15.6 0.284 0.930 64.5  99.6 4.0
22 2 23.02 −19 02.7 0.271 0.951 68.4  96.3 3.9
24 2 01.30 −16 26.3 0.260 0.972 72.9  92.3 4.0
26 1 38.70 −13 27.9 0.252 0.994 77.9  87.8 4.0
28 1 15.52 −10 11.2 0.246 1.017 83.3  82.8 4.0
30 0 52.10 − 6 42.0 0.244 1.041 89.0  77.5 4.1
2001 May 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green